OpenStudy (anonymous):

Multiply or divide as indicated. (3x)^12/(3x)^4

OpenStudy (anonymous):

Because you're dividing you can subtract your exponents.

OpenStudy (anonymous):

It would be x^8 because the 3's can cancel out and subtract the exponents.

OpenStudy (anonymous):

You can cancel the 3's if it's \[\frac{3x^{12}}{3x^4}\] As written it's \[\frac{(3x)^{12}}{(3x)^4}\] Which is totally different!

OpenStudy (anonymous):

That's how I read it, Sacapuntus. :) @magan , which one does your problem look like? The top one or the bottom one? My answer is for the bottom (which is how you wrote it). tn.n's answer is for the top.

OpenStudy (anonymous):

its like the bottom one

OpenStudy (anonymous):

\[\frac{n^x}{n^y} = n^{x-y}\]

OpenStudy (vjordillo):

I agree to @SACAPUNTAS . They're different. Because if you put 3 with x inside the parenthesis, it means that including 3 will carry the same exponent and x will have the exponent that 3 has.